OpenStudy (anonymous):

Solve using the square root method. Round to the nearest tenth, if necessary. x2 + 15 =0

OpenStudy (anonymous):

would this be 3.9, -3.9?

OpenStudy (anonymous):

Are you sure it isn't: x^2 -15=0

jimthompson5910 (jim_thompson5910):

No, subtract 15 from both sides to get x^2 = -15 Then take the square root of both sides to get x = sqrt(-15) or x = -sqrt(-15) Notice the negative number in the square root. This means that you won't get a real number for your answer.

OpenStudy (anonymous):

not its x^2 + 15 = 0

OpenStudy (anonymous):

Otherwise, you are correct!

OpenStudy (anonymous):

Then there is no real answer.

OpenStudy (anonymous):

alright . thanks!